Find the number that is 3/4 of the way from 2/5 to 4 2/3.
Answer by josmiceli(19441)

The number you need is
3/4 of the way between the given numbers
-----------------------------
Convert 2/5 to 6/15
Likewise, 4 2/3 = 4 10/15
4 10/15 minus 6/15 = 4 4/15
( this is the interval between 2/5 and 4 2/3 )
----------------------------
Now take 3/4 of the interval
1/4 x 4 4/15 = 1 1/15
3 x 1 1/15 = 3 3/15
3 3/15 = 3 1/5
--------------------------
Now add this to 2/5, ( 6/15 )
3 3/15 + 6/15 = 3 9/15
3 9/15 = 3 3/5 answer
----------------------
check:
If I take 1/3 of 3 1/5, then there should
be 4 of those lengths in the interval
1/3 x 3 1/5 = 1 1/15
and
4 x 1 1/15 = 4 4/15 ( which is the interval
between 2/5 and 4 2/3 )
----------------------
check my math and maybe get a 2nd opinion, too